# Examining the robustness of LLM evaluation to the distributional assumptions of benchmarks

### v1

@@ 1 Introduction
Due to their effectiveness, LLMs are becoming very popular in both academia and industry, making their evaluation crucial. However, this effectiveness comes at the cost of increased complexity, which makes their evaluation very challenging. Although prior research has introduced benchmarks for different tasks along with evaluation measures, these assessments often overlook potential biases. When a benchmark includes multiple prompts with similar characteristics, it can increase or decrease the average performance of a model, so model comparisons can become brittle with respect to benchmark composition. In this work, we show that the inherent connections between the prompts in current benchmarks impact the models’ performance and their relative rankings.
@@ 2 Related work
Evaluating the performance of LLMs has become a critical area of research, drawing significant attention in recent years. Comprehensive surveys of LLM evaluation can be found in Chang et al. (2023); Guo et al. (2023), and Liang et al. (2022).
When assessing the quality of LLMs, the robustness aspect is becoming of increasing importance (Wang et al., 2022; Goel et al., 2021). Robustness investigates the stability of a model when confronted with unforeseen prompts. Robustness research can be divided into three main lines of work (Li et al., 2023): (i) robustness under distribution shift (Wang et al., 2021; Yang et al., 2023), (ii) robustness to adversarial input (Zhu et al., 2023; Wang et al., 2023a), and (iii) robustness to dataset bias (Gururangan et al., 2018; Le Bras et al., 2020; Niven and Kao, 2019). Our work falls into the latter category.
@@
Limited research has been conducted to study inherent biases in LLM benchmarks. Among existing works, Gururangan et al. (2018) and Niven and Kao (2019) have shown that models leverage spurious statistical relationships in the benchmark datasets and, thus, their performance on the benchmarks is overestimated. In the same spirit, Le Bras et al. (2020) propose to investigate AFLITE (Sakaguchi et al., 2023), an iterative approach to filter datasets by removing biased data points to mitigate overestimation of language models’ performance. More recently, Alzahrani et al. (2024) show that performance of LLMs is highly sensitive to minor changes in benchmarks with multiple-choice questions.
@@ 5.1 Performance vectors are correlated
For CNN/Daily Mail, the performance matrix contains ROUGE scores, which are continuous values. Thus, we use cosine similarity to compare the average correlations obtained from the original and permuted performance matrices. The results show that the correlations among original performance vectors are significantly greater.
@@ 5.3 Relationship between model performance and semantic similarity of prompts
Hellaswag and CommonsenseQA use a multiple-choice format. The lack of strong evidence supporting the correlation in these benchmarks (see Figures 4(b) and 4(c)) is likely due to the embeddings picking up similarities between the different choices, rather than the logic the LLMs employ to arrive at their conclusion. This is consistent with our findings for ANLI, where a significant relationship does not stem from inputs to the model, but from the LLMs’ failure points.
@@ Appendix A Prompt structure
The prompts used for inference are depicted in Figures 567 and 8 for ANLI, HellaSwag, CommonsenseQA and CNN/Daily respectively.

### v2

@@ 1 Introduction
Due to their effectiveness, LLMs are becoming very popular in both academia and industry, making their evaluation crucial. However, this effectiveness comes at the cost of increased complexity, which makes their evaluation very challenging. Although prior research has introduced benchmarks for different tasks along with evaluation measures, these assessments often overlook potential biases. When a benchmark includes multiple prompts with similar characteristics, it can increase or decrease the average performance of a model, so model comparisons can become brittle with respect to benchmark composition (see Figure 1 for an illustrative example). In this work, we show that the inherent connections between the prompts in current benchmarks impact the models’ performance and their relative rankings.
@@
(a) All prompts contribute equally during evaluation.
(b) Prompts are weighted during evaluation.
Figure 1: Illustrative example showcasing how different distributional assumptions of benchmarks affect model rankings. Consider a benchmark containing prompts reflecting three different tasks: math (red triangles), code generation (blue circles), and text generation (green squares). In Figure 1(a), each benchmark prompt contributes equally to the model evaluation. In contrast, Figure 1(b) accounts for correlations between prompts and the weights of the prompts are adjusted accordingly during evaluation. In scenario 1(a), the red LLM ranks highest because it excels in math, and the benchmark is biased towards math tasks (7 out of 12 prompts are math-related). When considering different weights in scenario 1(b), we observe a different ranking outcome.
@@ 2 Related work
Evaluating the performance of LLMs has become a critical area of research, drawing significant attention in recent years. Comprehensive surveys of LLM evaluation and benchmark quality can be found in Chang et al. (2023); Guo et al. (2023); Perlitz et al. (2023), and Liang et al. (2022).
When assessing the quality of LLMs, the robustness aspect is becoming of increasing importance (Wang et al., 2022; Goel et al., 2021). Robustness investigates the stability of a model when confronted with unforeseen prompts. Robustness research can be divided into four main lines of work (Li et al., 2023): (i) robustness under distribution shift (Wang et al., 2021; Yang et al., 2023), (ii) robustness to adversarial input (Zhu et al., 2023; Wang et al., 2023a), (iii) robustness to prompt formats, including instruction templates (Mizrahi et al., 2023; Voronov et al., 2023; Weber et al., 2023; Sclar et al., 2023), and (iv) robustness to dataset bias (Gururangan et al., 2018; Le Bras et al., 2020; Niven and Kao, 2019). Our work falls into the latter category.
@@
Limited research has been conducted to study inherent biases in LLM benchmarks. Among existing works,  Gururangan et al. (2018) and Niven and Kao (2019) have shown that models leverage spurious statistical relationships in the benchmark datasets and, thus, their performance on the benchmarks is overestimated. In the same spirit, Le Bras et al. (2020) propose to investigate AFLITE (Sakaguchi et al., 2023), an iterative approach to filter datasets by removing biased data points to mitigate overestimation of language models’ performance. More recently, Alzahrani et al. (2024) show that performance of LLMs is highly sensitive to minor changes in benchmarks with multiple-choice questions. Other studies demonstrate that benchmarks often include redundancy and effective LLM evaluation can be achieved with a significantly smaller sample size (Polo et al., 2024; Vivek et al., 2024).
@@ 5.1 Performance vectors are correlated
For CNN/Daily Mail the performance matrix contains ROUGE scores, which are continuous values. Thus, we use cosine similarity to compare the average correlations obtained from the original and permuted performance matrices. The results show that the correlations among original performance vectors are significantly greater.
@@ 5.3 Relationship between model performance and semantic similarity of prompts
Hellaswag and CommonsenseQA use a multiple-choice format. The lack of strong evidence supporting the correlation in these benchmarks (see Figures 5(b) and 5(c)) is likely due to the embeddings picking up similarities between the different choices, rather than the logic the LLMs employ to arrive at their conclusion. This is consistent with our findings for ANLI, where a significant relationship does not stem from inputs to the model, but from the LLMs’ failure points.
@@ Appendix A Prompt structure
The prompts used for inference are depicted in Figures 678 and 9 for ANLI, HellaSwag, CommonsenseQA and CNN/Daily respectively.